OpenStudy (anonymous):

In the reaction, CH4 + 2O2 → CO2 + 2H2O, 8.9 × 105 joules of heat are released. Using Einstein’s equation, calculate the mass amount to which this energy equates. [(1 J = 1 kg times m squared over s squared) and the speed of light = 3.00 × 108 m over s] 9.9 ×1012 kg 9.9 × 10-12 kg 9.9 × 10-14 kg 1.1 × 10-6 kg @Abmon98 @Abhisar

OpenStudy (anonymous):

@Abmon98 I need you help please

OpenStudy (abmon98):

E=mc^2 8.9 × 10^5=m*(3.00*10^8)^2 m=8.9*10^5/(3.00*10^8)^2

OpenStudy (anonymous):

b right heh @Abmon98

OpenStudy (anonymous):

@Abmon98 I need you help on two more is that ok

OpenStudy (abmon98):

sure :)

OpenStudy (anonymous):

ok on here or make a new one @Abmon98

OpenStudy (abmon98):

post a new one

OpenStudy (anonymous):

ok :) @Abmon98